If $8500 is invested at a rate of 6% per year compounded daily, what is the balance after 4 years?
Answer by nerdybill(7384)

If $8500 is invested at a rate of 6% per year compounded daily, what is the balance after 4 years?
.
A = P(1+r/n)^(nt)
where
A is amount after time t (what we're looking for)
P is principal (8500)
r is rate (.06)
n is times compounded (365)
t is years invested (4)
.
A = P(1+r/n)^(nt)
A = 8500(1+.06/365)^(365*4)
A = 8500(1+.000164384)^(365*4)
A = 8500(1.000164384)^1460
A = 8500(1.271224077)
A = $10805.40